OpenStudy (anonymous):

A right triangle has a hypotenuse of 70 ft. and a leg of 35 ft., what is the length of the other leg?

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 35^2+b^2 = 70^2 1225+b^2 = 4900 b^2 = 4900 - 1225 b^2 = 3675 b = sqrt(3675) b = 60.6217783 So the other leg is roughly 60.6217783 ft

OpenStudy (anonymous):

let x be the unknown lenght so using pythangoras theorem x^2 + 35^2 = 70^2 so x^2 = 70^2 - 35^2 x^2 = 3675 and x = \[35\sqrt{5}\]